(1 + 1) REMPI SPECTROSCOPY OF NO.N_{2} and NO.CO VIA THE Ã STATE

\underline{TIMOTHY G. WRIGHT}, SOPHIA E. DAIRE, JÈROME LOZEILLE, STUART D. GAMBLIN, School of Chemistry, Physics and Environmental Sciences, University of Sussex, Falmer, Brighton, BN1 9QJ, U. K..

The Ã ← X̃ transition of the NO.N_{2} and NO.CO molecular complexes have been recorded using one-color (1 + 1) REMPI spectroscopy. The spectra exhibit vibronic features, which indicate some of the underlying rotational structure. Some insights into the structure of the spectra are gleaned from ab initio calculations.